Find 2 numbers whose sum is 92, if the first is 4 more than 7 times the second?
Answer by Earlsdon(6294)

Let x = the first number, and the second number y.
x = 7y+4 and...
x+y = 92 Substitute for x: x = 7y+4
(7y+4)+y = 92 Solve for y.
8y+4 = 92
8y = 88
y = 11 and...
x = 7y+4
x = 77+4
x = 81
--------------
The two numbers are: 81 and 11
---------------